OpenStudy (anonymous):

a point on the terminal side of angle theta is given. find the exact value ofthe indicated trigonometric function of theta. (1,-4), find theta.

Miracrown (miracrown):

so, let's see where the point is, first ...

OpenStudy (anonymous):

i got tan theta= y/x= -4/1=-4 is this correct?

Miracrown (miracrown):



Miracrown (miracrown):

so there is the point ^

OpenStudy (anonymous):

is my solution correct?

OpenStudy (anonymous):

look i typed a solution

Miracrown (miracrown):

yes, correct :)

OpenStudy (anonymous):

hey i have a question

Miracrown (miracrown):

Sure, what is it?

Miracrown (miracrown):

so, we need to find the angle in the fourth quadrant that has tan(that angle) = -4

OpenStudy (anonymous):

why? i thought you said the solution was correct

OpenStudy (anonymous):

but if you want the theta in q4, theta=arctan-4 =-76 degrees.

Miracrown (miracrown):

so, the angle we want will be tan^-1(-4) that's not a nice angle, so since they want it exact, they likely want us to leave it in that form

OpenStudy (anonymous):

multiple choice has -4,4,1/4,4/1 for answers. we simplified enough

Miracrown (miracrown):

approximately if you don't want to use negative angles, you could use 360+tan^-1(-4)

OpenStudy (anonymous):

ok, ready for my question?

Miracrown (miracrown):

wait, the question says indicated trigonometric function of theta what function do they want us to find?

OpenStudy (anonymous):

i typed it word for word. i didnt leave anything out. im sure this is right due to -4 being one of the answers

Miracrown (miracrown):

yes, it's likely they wanted tangent though leaving that out is probably a typo :)

OpenStudy (anonymous):

yes tan theta, srry

OpenStudy (anonymous):

ready for my other question?

OpenStudy (anonymous):

im going make a new post so u will earn easy medal